Task 1: Organization  Task 1: Organization MFG discussion document  Task 1: Organization  1

Note on math font group organization

Alan Jeffrey

This note is based on an electronic mail discussion between Alan Jeffrey and Justin Ziegler. Our suggested breakdown of tasks for the math font group (MFG) is:

1. Organization: writing papers such as this one, about the structure and direction of the group.

2. Requirements analysis: looking at the needs and current practice of mathematical typesetting, and setting (perhaps unachievable!) goals.

3. Technical studies: individual technical analysis, with clear conclusions and stating the advantages and disadvantages of the approaches investigated.

4. Standardization: recording the WG’s decisions, leading towards the proposal of new math font encodings and related standards.

5. Implementation: designing software to support the new encodings.

These are only rough suggestions, all better suggestions welcomed! We might want to break each task down into subtasks. Organization could include:

• Statement of purpose: a one-page document setting out the WG’s aims.

• Outlining goals: general papers like this one.

• Setting timetables: completion dates for these

goals. Requirements analysis could include:

• Common mathematical usage: what features are common across different subjects?

• Specific mathematical usage: what features are specific to particular subjects? (For example setting of scripts on roman letters in chemistry, multiletter identifiers in computer science…)

• Non-mathematical usage: what is math mode used for other than for mathematics? (For example, setting computer programs, fractions, footnote markers, trademark signs…)

• Compatability: how compatible should a new encoding be with Knuth’s?

• Extensions: what new symbols and other features are users crying out for?

• Non-CM fonts: what are the requirements for users of non-CM fonts (for example PostScript outline fonts).

Technical studies could include:

• The current state of TEX : what are the good and bad points of TEX’s current mathematical typesetting?

• The limitations of TEX : what restrictions does TEX make to mathematical typesetting? (For example, the math spacing table in Chapter 18 of The TEXbook is hard-wired, Appendix G of The TEXbook overloads many font dimensions…)

• Prototype implementations: each part of the standard should be α-tested before standardization.

• Studying requirements: work on problems suggested by the requirements analysis task.

Standardization could include:

• Font encodings: this is probably the most important point!

• Encoding vectors: naming all the symbols as PostScript encoding vectors.

• Font information: any new font dimensions, kerning information, charlists, and so on.

• (L^{A})TEX interface: how the fonts can be used in (L^{A})TEX, for example giving default control sequence names for each glyph.

Implementation could include:

• programs: implementing of the encodings using character shapes based on cmmi,

                                              cmsy and cmex.

• TEX programs: a replacement for plain.tex, and document styles for using the font encodings with L^{A}TEX.

• VPL manipulation tools: to help install Non-CM fonts, we may need general virtual font manipulators.

• Benchmarks: equivalents of the trip.tex and testfont.tex tests for the new encodings.

⋄ Alan Jeffrey University of Sussex

                                                                alanje@cogs.susx.ac.uk